OpenStudy (andijo76):

A bag contains 6 purple marbles, 7 azure marbles and 10 orange marbles. What is the chance of drawing an orange marble? If an orange marble is drawn then placed back into the bag, and a second marble is drawn, what is the probability of drawing an azure marble? Give solutions exactly in reduced fraction form, separated by a comma

OpenStudy (akshay_budhkar):

Total number of marbles = 23. Now we have \[P(orange)= {No. of orange \over total no.}={ 10 \over 23}\]

OpenStudy (lgbasallote):

probability of orange marble = 10/23 probability of azure = 7/23

OpenStudy (akshay_budhkar):

Similarly for azure marble you have \[7 \over 23\]